# Phase-field modeling of cyclic behavior in quasi-brittle materials: a micromechanics-based approach

### v2

@@ Abstract
In this paper, we extend the micromechanics-based phase-field modeling of fatigue fracture to capture cyclic plasticity with ratcheting. This mechanism is particularly important for low-cycle fatigue, where the accumulation of inelastic strains plays an important role in the progression to final failure. The ratcheting contribution is formulated through the evolution of ratcheting strain, which accumulates over loading cycles and captures the inelastic strain growth characteristic of cyclic plasticity in a thermodynamically consistent manner. The extended plastic potential allows independent control over deviatoric and volumetric ratcheting components. Numerical simulations are performed to evaluate the model under both monotonic and cyclic loading and to assess the influence of ratcheting on material response.
@@ 1 Introduction
The cyclic behavior of quasi-brittle materials has been addressed through two key components–fatigue and ratcheting–considering that both damage and plasticity exhibit different characteristics under cyclic loading. Fatigue refers to the progressive deterioration of a material under cyclic loading due to damage. Fatigue can occur due to repeated loading and unloading cycles, leading to microcrack initiation and propagation, ultimately resulting in failure Kalina [2024]. The understanding of fatigue behavior is challenging because of the complex interplay between microstructure, fatigue damage mechanisms, and loading conditions. Predicting fatigue behavior in quasi-brittle materials has been challenging and remains a topic of interest. Many studies have investigated fatigue phenomena, employing different experimental techniques such as acoustic emission and digital image correlation Deresse et al. [2022, 2023, 2024]. Moreover, various numerical methodologies and phenomenological formulations have been utilized to simulate this behavior. Among these methods, the phase-field approach to fracture is of particular interest due to the advantages discussed in the previous sections.
Different approaches incorporating fatigue into the phase-field model have been suggested in recent years. These phenomenological models aim to simulate fatigue behavior by introducing history-dependent functions that gradually diminish the material’s stiffness, ultimately leading to failure. Different models can be categorized based on their definition of the fatigue variable or how this variable is incorporated into the model to represent the material’s loading history Kalina et al. [2023]. Some methods include increasing the crack driving force by an additional energy term Schreiber et al. [2020], Haveroth et al. [2020], Loew et al. [2020], or reducing the fracture energy to simulate the diminishing resistance of the material against crack propagation, achieved through the implementation of a fatigue degradation function Alessi et al. [2018], Carrara et al. [2020], Seiler et al. [2020], Kristensen and Martínez-Pañeda [2020], Grossman-Ponemon et al. [2022]. In the first type of models, the additive fatigue contribution is an energetic quantity, whereas in the second type of models, the fatigue degradation function is scalar and without unit.
In the second type of fatigue models, the degradation function depends on a fatigue history variable, which evolves over time and is typically expressed in rate form due to its process-dependent nature. The history variable can be configured to increase only during unloading to avoid impacting the monotonic loading case, as demonstrated in Seleš et al. [2021], or only during loading, as shown in Carrara et al. [2020]. This variable accumulates over loading cycles, and once a specific threshold is surpassed, fatigue degradation begins to reduce the fracture energy. This model is originally developed for brittle fracture and has been further extended to ductile fracture Ulloa et al. [2021a], Khalil et al. [2022].
@@ 3 Incorporation of ratcheting-driven cyclic plasticity
A micromechanics-based phase-field model is proposed to capture both ratcheting-driven cyclic plasticity and fatigue. The formulation builds on the energetic framework of the micromechanics-based phase-field model introduced in You et al. [2021], Ulloa et al. [2022] and subsequently extended to fatigue in our earlier work Sarem et al. [2025]. In this framework, the phase-field method characterizes damage dissipation at the macro scale, while Drucker-Prager plasticity governs the evolution of plastic strain.
@@ 3.1 Free energy
Equation (2) shows that in the presence of open microcracks, frictional sliding is not expected, resulting in the exclusion of plastic strain from contributing to the strain energy density. Consequently, the model behaves in a brittle manner. On the other hand, when closed microcracks undergo sliding, both the stored energy and the blocked friction energy contribute to the strain energy density. The ratcheting strain appears in the elastic part of the free energy because it represents a slowly evolving, irreversible shift of the elastic reference state caused by microstructural rearrangements. In the rheological model (figure 1), it acts as a slider in series with the elastic spring, altering the stored elastic energy at each cycle.
@@
The kinematic hardening, as shown in equation (8), is obtained from the stiffness tensor and is linked to the degradation function $g(\alpha)$. Consequently, there is no longer a necessity for phenomenological definitions of kinematic hardening. $\bm{\mathsf{H}}^{\mathrm{kin}}(\alpha)$, which here couples damage and plasticity, approaches infinity as the damage variable tends to zero. To prevent this, an initial damage is always required, representing the existence of microcracks which aligns with the physical assumptions of the micromechanics-based framework. The use of a small initial value has no effect on the final damage profile, as discussed in Ulloa et al. [2022].
@@
In the present formulation, the introduction of the ratcheting strain $\bm{\varepsilon}^{\mathrm{r}}$ results in a controlled, regime-switching discontinuity in the free energy. The switching condition between open and closed microcracks (see equation (13)) ensures that this discontinuity occurs only at physically meaningful transitions, corresponding to the onset or loss of contact between crack faces. This discontinuity is therefore a representation of two distinct micro-mechanical states: open microcracks, where no contact or friction occurs, and closed microcracks, where sliding and ratcheting mechanisms are activated.
The following sections show that dissipation is consistently formulated through convex, non-negative potentials, ensuring that the overall formulation remains thermodynamically admissible. Consequently, the Clausius–Duhem inequality is satisfied even in the presence of non-smooth energy branches. The accumulated ratcheting strain introduces an additional irreversible, dissipative mechanism without violating the variational structure of the model Francfort [2018].
@@ 3.2 Plasticity criterion with a non-associated plastic flow rule
To incorporate ratcheting behavior into the flow rule, the plastic potential function $g^{\mathrm{p}}$ is defined as follows Ulloa et al. [2021b]:
@@
where $A_{\theta}$ is the dilation coefficient which is assumed to take a value between zero and $A_{\varphi}$. $\beta_{K}$ and $\beta_{\mu}$ are ratcheting parameters associated with volumetric and deviatoric contributions, respectively. Since the underlying plasticity model is pressure-dependent, the ratcheting mechanism is also formulated to account for both volumetric and deviatoric strain components Ulloa et al. [2021b]. In this context, macroscopic dilatancy is directly related to the normal opening of microcracks during frictional sliding, caused by the roughness of these microcracks. It is important to note that despite the non-associative nature of the plasticity model, it remains variational. Further details on this can be found in Ulloa et al. [2021b, 2022].
@@ 3.4 Dissipation potential
The dissipation potential characterizes the energy dissipated during irreversible processes such as damage, plastic deformation and ratcheting. It is defined as a function of the rates of internal variables. A thermodynamically admissible formulation of the dissipation potential is expressed as the sum of plastic and damage contributions Alessi et al. [2019], Ulloa et al. [2021a]:
@@
To define the plastic dissipation, the set of admissible stress states must first be specified. As shown in Ulloa et al. [2021b], the non-associative model maintains a variational structure if a state-dependent convex set is defined as follows Francfort [2018]:
@@
This extended formulation effectively captures the progressive and irreversible nature of ratcheting in materials subjected to cyclic loading. The explicit control provided by the ratcheting parameters $\beta_{\mu}$ and $\beta_{K}$ enables the simulation of complex cyclic behavior, such as mean strain accumulation and asymmetric hysteresis loops.
@@ 4 Numerical implementation
Algorithm 1 provides an overview of the staggered procedure employed in this study. To begin with, the nonlinear mechanical equilibrium equation is solved to find $\bm{u}$. Subsequently, the damage evolution problem is solved to obtain $\alpha$ using the history-field approach Miehe et al. [2010b]. Next, a return-mapping algorithm is employed to update $\bm{\varepsilon}^{\mathrm{p}}$. In this sub-problem, the algorithm determines the state of the solid matrix, denoted as state, in terms of open or closed microcracks using equation (13). Through this iterative procedure, the governing equations are solved at each loading step and for every Gauss point, ensuring a consistent update of the field variables.
At each iteration, one field variable is solved while keeping the others fixed. The updated field is then incorporated into the system, and the process is repeated for the remaining variables until convergence is achieved. This sequential update ensures numerical stability while effectively handling the strong coupling between mechanical equilibrium, plasticity, and phase-field damage evolution. The detailed solution procedure for the three sub-problems is provided in the next subsections.
@@ 4.1 Mechanical balance equation
First, the equilibrium problem is solved for the displacement $\bm{u}$ by linearizing the mechanical balance equation, while the other two variables $\bm{\varepsilon}^{\mathrm{p}}$ and $\alpha$ are fixed. The weak form of the mechanical balance is given by
@@ 4.2 Plastic evolution problem
    $=\bm{\mathsf{C}}-\frac{\big(2\mu\,\hat{\bm{n}}_{\mathrm{dev}}^{\mathrm{trial}}+\sqrt{6}\,KA_{\theta}\bm{1}\big)\otimes\big(2\mu\,\hat{\bm{n}}_{\mathrm{dev}}^{\mathrm{trial}}+\sqrt{6}\,KA_{\varphi}\bm{1}\big)}{[2\mu+H^{\mathrm{kin}}_{\mu}(\alpha)]+6A_{\varphi}A_{\theta}\big[K+H^{\mathrm{kin}}_{K}(\alpha)\big]}-\frac{4\mu^{2}\Delta\gamma}{\|\bm{s}^{\mathrm{p\,trial}}_{\mathrm{dev}}\|}\bigg(\bm{\mathsf{I}}-\frac{1}{3}\bm{1}\otimes\bm{1}-\hat{\bm{n}}_{\mathrm{dev}}^{\mathrm{trial}}\otimes\hat{\bm{n}}_{\mathrm{dev}}^{\mathrm{trial}}\bigg).$
@@
4.3 Damage evolution problem
The equilibrium and plastic evolution problems are solved in the previous steps to obtain the displacements $\bm{u}$ and plastic strains $\bm{\varepsilon}^{\mathrm{p}}$. The subsequent step in the staggered solution approach is to address the following variational inequality to obtain $\alpha$
$\Delta\alpha\geq 0.$ (65)
[…]
and $s^{\mathrm{d\,hist}}$ is defined as
[…]
Subsequently, equation (64) is written as
@@ 5.3 Ratcheting response of a material point
Figure 15: Response of the material point under axial strain-controlled loading: (a) stress-strain curve, (b) evolution of damage () and fatigue degradation function (), (c) evolution of the deviatoric stress component $\sigma_{\mathrm{dev}\,xx}$, and (d) evolution of axial ratcheting strain $\bm{\varepsilon}^{\mathrm{r}}_{xx}$ for two parameter sets: $\beta_{K}=0.5$, $\beta_{\mu}=0.4$ (), and $\beta_{K}=0.2$, $\beta_{\mu}=0.1$ ().
@@ 5.4 Cyclic response of a perforated specimen
Figure 17: (a) Force-displacement response of the perforated specimen under cyclic loading, and (b) evolution of axial ratcheting strain $\bm{\varepsilon}^{\mathrm{r}}_{yy}$ at point A for two parameter sets: $\beta_{K}=0.4$, $\beta_{\mu}=0.3$ (), and $\beta_{K}=0.2$, $\beta_{\mu}=0.1$ ().

### v3

@@ Abstract
In this paper, we extend a micromechanics-based phase-field framework for fatigue fracture to incorporate cyclic plasticity with ratcheting. This mechanism is particularly relevant for low-cycle fatigue, where the accumulation of inelastic strains plays a critical role in the progression to final failure. An energetic formulation is proposed in which the ratcheting strain is explicitly incorporated into both the free energy and the dissipation potential. Ratcheting is modeled within a pressure-dependent, non-associative plasticity framework through the evolution of a ratcheting strain that progressively accumulates over loading cycles, capturing the characteristic inelastic strain growth of cyclic plasticity in a thermodynamically consistent manner. The plastic potential is formulated such that the deviatoric and volumetric components of ratcheting can be controlled independently. A staggered solution scheme is employed to solve for the internal variables, including the ratcheting strain. Numerical examples under monotonic and cyclic loading conditions are presented to evaluate the proposed model and to investigate the influence of ratcheting on the material response.
@@ 1 Introduction
The cyclic behavior of quasi-brittle materials has been addressed through two key components–fatigue and ratcheting–considering that both damage and plasticity exhibit different characteristics under cyclic loading. Fatigue arises from repeated loading and unloading cycles, leading to microcrack initiation and propagation and ultimately resulting in failure Kalina [2024]. Understanding fatigue behavior remains challenging due to the complex interplay between material microstructure, fatigue damage mechanisms, and loading conditions. Numerous experimental studies have investigated fatigue phenomena using techniques such as acoustic emission and digital image correlation Deresse et al. [2022, 2023, 2024]. In parallel, various numerical methodologies and phenomenological formulations have been developed to simulate fatigue behavior. These phenomenological models typically rely on history-dependent functions that progressively degrade the material stiffness, ultimately leading to failure.
Phenomenological fatigue models can be broadly categorized based on the definition of the fatigue variable and the manner in which it is incorporated to represent the material’s loading history Kalina et al. [2023]. One class of models introduces fatigue effects through an additional energetic contribution that increases the crack driving force Schreiber et al. [2020], Haveroth et al. [2020], Loew et al. [2020]. Another class accounts for fatigue by reducing the fracture energy via a scalar fatigue degradation function, thereby representing the progressive loss of resistance to crack propagation Alessi et al. [2018], Carrara et al. [2020], Seiler et al. [2020], Kristensen and Martínez-Pañeda [2020], Grossman-Ponemon et al. [2022]. In the former approach, the fatigue contribution enters as an energetic quantity, whereas in the latter, the degradation function is dimensionless.
In degradation-based fatigue models, the degradation function depends on a fatigue history variable that evolves over time and is typically formulated in rate form due to its process-dependent nature. This history variable may be designed to evolve only during unloading, in order to preserve the monotonic response Seleš et al. [2021], or only during loading Carrara et al. [2020]. The variable accumulates over loading cycles, and once a prescribed threshold is exceeded, fatigue degradation initiates and progressively reduces the fracture energy. Although originally developed for brittle fracture, this modeling framework has subsequently been extended to ductile fracture Ulloa et al. [2021a], Khalil et al. [2022].
@@ 3 Incorporation of ratcheting-driven cyclic plasticity
A micromechanics-based phase-field model is proposed to capture both ratcheting-driven cyclic plasticity and fatigue. The formulation builds upon the energetic framework of the micromechanics-based phase-field model introduced in You et al. [2021], Ulloa et al. [2022] and subsequently extended to fatigue in our earlier work Sarem et al. [2025]. Within this framework, damage dissipation at the macroscopic scale is described by the phase-field approach, while the evolution of plastic strain is governed by Drucker–Prager plasticity. The present section details the extension of the model to account for ratcheting effects through the incorporation of ratcheting strain into both the free energy and the dissipation potential in a thermodynamically consistent manner.
@@ 3.1 Free energy
Equation (2) shows that in the presence of open microcracks, frictional sliding is not expected, resulting in the exclusion of plastic strain from contributing to the strain energy density. Consequently, the model behaves in a brittle manner. On the other hand, when closed microcracks undergo sliding, both the stored energy and the blocked friction energy contribute to the strain energy density. The ratcheting strain is included in the elastic part of the free energy because it represents a slowly evolving, irreversible shift of the elastic reference state caused by microstructural rearrangements. In the rheological model (figure 1), it acts as a slider in series with the elastic spring, altering the stored elastic energy at each cycle.
@@
The kinematic hardening, as shown in equation (8), is obtained from the stiffness tensor and is linked to the degradation function $g(\alpha)$. Consequently, there is no longer a necessity for phenomenological definitions of kinematic hardening. $\bm{\mathsf{H}}^{\mathrm{kin}}(\alpha)$, which here couples damage and plasticity, approaches infinity as the damage variable tends to zero. To prevent this, an initial damage is always required, representing the existence of microcracks which aligns with the physical assumptions of the micromechanics-based framework. The use of a small initial value has no effect on the final damage profile Ulloa et al. [2022].
@@
In the present formulation, the introduction of the ratcheting strain $\bm{\varepsilon}^{\mathrm{r}}$ results in a controlled, regime-switching discontinuity in the free energy. The switching condition between open and closed microcracks (see equation (13)) ensures that this discontinuity occurs only at physically meaningful transitions, corresponding to the onset or loss of contact between crack faces. This discontinuity is therefore a representation of two distinct micromechanical states: open microcracks, where no contact or friction occurs, and closed microcracks, where sliding and ratcheting mechanisms are activated.
The following subsections show that dissipation is consistently formulated through convex, non-negative potentials, ensuring that the overall formulation remains thermodynamically admissible. Consequently, the Clausius–Duhem inequality is satisfied even in the presence of non-smooth energy branches. The accumulated ratcheting strain introduces an additional irreversible, dissipative mechanism without violating the variational structure of the model Francfort [2018].
@@ 3.2 Plasticity criterion with a non-associated plastic flow rule
To incorporate ratcheting behavior into the flow rule, the plastic potential function $g^{\mathrm{p}}$ is defined as follows:
@@
where $A_{\theta}$ is the dilation coefficient which is assumed to take a value between zero and $A_{\varphi}$. $\beta_{K}$ and $\beta_{\mu}$ are ratcheting parameters associated with volumetric and deviatoric contributions, respectively. Since the underlying plasticity model is pressure-dependent, the ratcheting mechanism is also formulated to account for both volumetric and deviatoric strain components. In this context, macroscopic dilatancy is directly related to the normal opening of microcracks during frictional sliding, caused by the roughness of these microcracks. It is important to note that despite the non-associative nature of the plasticity model, it remains variational. For further details, the reader is referred to Ulloa et al. [2021b].
@@ 3.4 Dissipation potential
The dissipation potential characterizes the energy dissipated during irreversible processes such as damage, plastic deformation and ratcheting. It is defined as a function of the rates of internal variables. A thermodynamically admissible formulation of the dissipation potential is expressed as the sum of plastic and damage contributions Alessi et al. [2019]:
@@
To define the plastic dissipation and incorporate ratcheting effects, the set of admissible stress states is first specified. The non-associative model maintains a variational structure if a state-dependent convex set is defined as follows Francfort [2018]:
@@
The derived extended plastic dissipation potential effectively captures the progressive and irreversible nature of ratcheting under cyclic loading. The explicit control provided by the ratcheting parameters $\beta_{\mu}$ and $\beta_{K}$ enables the simulation of complex cyclic behavior, such as mean strain accumulation and asymmetric hysteresis loops.
@@ 4 Numerical implementation
Algorithm 1 provides an overview of the staggered procedure employed in this study. To begin with, the nonlinear mechanical equilibrium equation is solved to find $\bm{u}$. Subsequently, the damage evolution problem is solved to obtain $\alpha$ using the history-field approach Miehe et al. [2010b]. Next, a return-mapping algorithm is employed to update $\bm{\varepsilon}^{\mathrm{p}}$ and $\bm{\varepsilon}^{\mathrm{r}}$. In this sub-problem, the algorithm determines the state of the solid matrix, denoted as state, in terms of open or closed microcracks using equation (13). Through this iterative procedure, the governing equations are solved at each loading step and for every Gauss point, ensuring a consistent update of the field variables. The detailed solution procedure for the three sub-problems is provided in the next subsections.
@@ 4.1 Mechanical balance equation
First, the equilibrium problem is solved for the displacement $\bm{u}$ by linearizing the mechanical balance equation, while the other three variables $\bm{\varepsilon}^{\mathrm{p}}$, $\bm{\varepsilon}^{\mathrm{r}}$ and $\alpha$ are fixed. The weak form of the mechanical balance is given by
@@
4.2 Damage evolution problem
The equilibrium is solved in the previous step to obtain the displacement $\bm{u}$. The subsequent step in the staggered solution approach is to address the following variational inequality to obtain $\alpha$
$\Delta\alpha\geq 0.$ (50)
[…]
and $s^{\mathrm{d\,hist}}$ is defined as
[…]
Subsequently, equation (49) is written as
@@ 4.3 Plastic evolution problem
    $=\bm{\mathsf{C}}-\frac{\big(2\mu\,\hat{\bm{n}}_{\mathrm{dev}}^{\mathrm{trial}}+\sqrt{6}\,KA_{\theta}\bm{1}\big)\otimes\big(2\mu\,\hat{\bm{n}}_{\mathrm{dev}}^{\mathrm{trial}}+\sqrt{6}\,KA_{\varphi}\bm{1}\big)}{[2\mu+H^{\mathrm{kin}}_{\mu}(\alpha)]+6A_{\varphi}A_{\theta}\big[K+H^{\mathrm{kin}}_{K}(\alpha)\big]}$
    $\hskip 62.59596pt-\frac{4\mu^{2}\Delta\gamma}{\|\bm{s}^{\mathrm{p\,trial}}_{\mathrm{dev}}\|}\bigg(\bm{\mathsf{I}}-\frac{1}{3}\bm{1}\otimes\bm{1}-\hat{\bm{n}}_{\mathrm{dev}}^{\mathrm{trial}}\otimes\hat{\bm{n}}_{\mathrm{dev}}^{\mathrm{trial}}\bigg).$
@@ 5.3 Ratcheting response of a material point
Figure 15: Response of the material point under axial strain-controlled loading: (a) stress-strain curve, (b) evolution of damage () and fatigue degradation function (), (c) evolution of the deviatoric stress component $\sigma_{\mathrm{dev}\,xx}$, and (d) evolution of axial ratcheting strain $\bm{\varepsilon}^{\mathrm{r}}_{xx}$ for two parameter sets: $\beta_{K}=0.2$, $\beta_{\mu}=0.1$ (), and $\beta_{K}=0.5$, $\beta_{\mu}=0.4$ ().
@@ 5.4 Cyclic response of a perforated specimen
Figure 17: (a) Force-displacement response of the perforated specimen under cyclic loading, and (b) evolution of axial ratcheting strain $\bm{\varepsilon}^{\mathrm{r}}_{yy}$ at point A for two parameter sets: $\beta_{K}=0.2$, $\beta_{\mu}=0.1$ (), and $\beta_{K}=0.4$, $\beta_{\mu}=0.3$ ().